# UPose3D: Uncertainty-Aware 3D Human Pose Estimation with Cross-View and Temporal Cues

### v2

@@ V-D Ablation Study
Inputs of Pose Compiler. Finally, we investigate the effect of different point cloud formation strategies in our pipeline. Specifically, we aim to study the impact of appending a relative camera position embedding, inspired by [6], to the cross-projected 2D keypoints while creating the point clouds. Accordingly, in our first experiment, we concatenate the epipolar line parameters of other views to the point cloud of the reference view. Similarly, in our second experiment, we concatenate the relative position of the other cameras to the input point cloud as well. However, as shown in Table VI, adding extra inputs does not have a large impact on the performance.
TABLE VII: Model parameters.
Method Parameters M)
Learnable Triangulation [2] 81
Epipolar Transformers [3] 69
MTF-Transformers [6] 72
AdaFuse [4] 69
Ours 65
V-E Computation
Table VI shows the inference time, computational cost, and the number of parameters of our network. We observe that our pose compiler is significantly smaller than a single 2D pose estimator, taking less than 1% of the total parameter count. Additionally, in Table VII, we present the number of parameters of our method in comparison to state-of-the-art 3D pose estimation methods. We show that our method is considerably lighter than prior works.
V-F Viewpoint Scalability
To test the scalability of our method with viewpoints, we evaluate it with different numbers of cameras on the CMU Panoptic [47] dataset and present the results in Fig. 5. For this experiment, we use the HRNet-W48 [24] backbone similar to our OoD experiments to analyze scalability in in-the-wild scenarios. Additionally, we report the error and runtime of two triangulation techniques, RANSAC and DLT, for comparison. We tune the RANSAC implementation of [2] and run it for 20 iterations. The left figure shows that as the number of views increases, the error of UPose3D continuously improves while DLT and RANSAC reach a plateau at around 85 mm. In the right figure, we show that the inference time of our pipeline with a single batch size remains less than RANSAC. However, increasing the batch size causes both improvements and variability in the runtime at the cost of more computation memory during MLE.
Figure 5: We demonstrate the scalability of UPose3D to the number of cameras and its Runtime.
Figure 6: Illustration of UPose3D on Human3.6m [11] (right) and RICH [12] (left) datasets, showing the accurate 3D pose estimated by our UPose3D (top) compared to ground-truth (bottom). The output keypoint distributions from our 2D pose estimator and pose compiler are also demonstrated as heatmaps.
V-G Qualitative Analysis
Fig. 6 depicts a few challenging examples and shows the outputs of our method. Specifically, by sampling it in every pixel, we demonstrate the output likelihood of keypoints from the 2D pose estimator and our pose compiler as heatmaps. Despite choosing challenging scenes, we observe that the predictions of our method are still close to the ground-truth keypoints, indicating that our method does not produce any significant outliers within its output even on the unseen samples of RICH [12] dataset. Additionally, we observe that the 2D pose estimator generally results in more refined predictions and sharper uncertainty distributions, while our pose compiler outputs a coarser distribution. Moreover, our method typically depicts higher horizontal uncertainties, which may be due to the human motions across several frames.
@@ V-G Qualitative Analysis
In Fig. 7, Fig. 8, and Fig. 9, we demonstrate more examples of our UPose3D on Human3.6m [11] dataset to showcase its visual fidelity in comparison to ground-truth keypoints and AdaFuse [4] during InD evaluation scheme. Additionally, we provide several visual examples of UPose3D results in Fig. 10, Fig. 11, and Fig. 12 in comparison to our implementation of AdaFuse [4] in OoD settings on the RICH [12] dataset. Unlike Fig. 6, to better visualize the sharp keypoint distribution output of our 2D pose estimators, we show the logarithm of heatmaps in all figures for the 2D pose estimators. We observe that Our method performs consistently in both settings, while AdaFuse fails to correctly predict the human keypoints in some OoD samples.
VI Conclusion
This paper presents UPose3D, a multi-view 3D human pose estimation method designed to address the challenges in generalization, scalability, and over-reliance on real-world 3D annotated data. It includes a novel cross-view fusion strategy that scales well with varying camera numbers and volume sizes. Additionally, UPose3D integrates a pose compiler that learns to predict keypoint positions and uncertainties given the cross-view and temporal information. This module is trained using our online multi-view data synthesis strategy. Finally, the uncertainties and predicted key points from two sources of image branch and pose compiler are used in an uncertainty-aware 3D human pose estimation algorithm. UPose3D outperforms state-of-the-art approaches in OoD evaluations while achieving competitive results to in-distribution methods without requiring annotated 3D pose data. In conclusion, UPose3D may positively impact applications in sports analysis and health care, where performance and robustness are crucial.
Limitations and Future Work. Our primary focus in this paper was the fidelity of estimated output 3D poses from multi-view inputs without requiring 3D supervision. We acknowledge, however, that our method could be optimized for real-time applications. A promising avenue to achieve faster inference involves the exploration of specialized second-order optimizers that use a deep-learning neural network to estimate the Hessian matrix during likelihood maximization. Furthermore, given the importance of uncertainty modeling in our performance, it may be possible to reach better robustness by introducing additional sources of constraints. For example, a depth estimation network in each view, as well as a trajectory prediction network, can be used as an additional term in the likelihood function to mitigate the effect of outliers and noisy predictions in 2D pose estimators.
Acknowledgment
This work was partially funded by Mitacs through the Accelerate program.

### v3

@@ 0.B.2 Additional Ablation Study
Inputs of Pose Compiler. Finally, we investigate the effect of different point cloud formation strategies in our pipeline. Specifically, we study the impact of appending a relative camera position embedding, inspired by [55], to the cross-projected 2D keypoints while creating the point clouds. Accordingly, in our first experiment, we concatenate the epipolar line parameters of other views to the point cloud of the reference view. Similarly, in our second experiment, we concatenate the relative position of the other cameras to the input point cloud as well. However, as shown in Tab. 5, adding extra inputs does not greatly impact the performance.
0.B.3 Additional Baselines and Comparisons
Table 6 complements Tab. 1 of the main paper by providing more comparisons with prior works on the Human3.6m [18] dataset. Here, we include 3D keypoint estimation approaches regardless of their input modality or supervision type in InD settings. We observe that our method outperforms all of the other approaches despite only using 2D supervision. Additionally, in Tab. 7, we compare our work with prior research on the RICH [16] dataset. Since this dataset was recently published, only monocular 3D body modeling techniques have reported their performance on this dataset. Here, we observe that our method outperforms the majority of prior works. More importantly, when comparing Tab. 6 and Tab. 7 we notice that our method achieves consistent results between InD and OoD evaluations on the Human3.6m [18] and the RICH [16] datasets, showing generalizability across in-studio and outdoor environments.
Appendix 0.C Additional Visual Examples
We provide a supplementary video that describes our method with visual demonstrations. Additionally, we provide several video clips of input and output data from Human3.6m [18], RICH [16], and CMU-Panoptic [22] datasets and compare the visual fidelity of our approach with the state-of-the-art method on Human3.6m, AdaFuse [66].
Appendix 0.D Qualitataive Comparisons
In Fig. 7, Fig. 8, and Fig. 9, we demonstrate some examples of our UPose3D on Human3.6m [18] dataset to showcase its visual fidelity in comparison to ground-truth keypoints and AdaFuse [66] in InD evaluation scheme. Additionally, we provide more visual examples of UPose3D results in Fig. 10, Fig. 11, and Fig. 12 in comparison to our implementation of AdaFuse [66] in OoD settings on the RICH [16] dataset. To better visualize the sharp keypoint distribution output of our 2D pose estimators, we show the logarithm of heatmaps in all figures for 2D pose estimators. We refer the reader to Fig. 4 of the main paper for an illustration of the real heatmaps without any post-processing. Our method performs consistently in both settings, while AdaFuse fails to correctly predict the human keypoints in some OoD samples. In all cases, the 2D pose estimator generally results in more refined predictions and sharper uncertainty distributions, while our pose compiler outputs a coarser distribution. Moreover, our method typically depicts higher horizontal uncertainties, which may be due to more frequent horizontal movements.